# “Sometimes You Just Gotta Risk It for the Biscuit [Uncaptioned image]”: A Portrait of Student Risk-Taking

### v2

@@ Abstract.
Understanding how individuals, including students, make decisions involving risk is a fundamental aspect of behavioral research. Despite the ubiquity of risk in various aspects of life, limited empirical work has explored student risk-taking behavior in computing education. This study aims to partially replicate prior research on risk-taking behavior in software engineers while focusing on students, shedding light on the factors that affect their risk-taking choices. In our work, students were presented with a hypothetical scenario related to meeting a course project deadline, where they had to choose between a risky option and a safer alternative. We examined several factors that might influence these choices, including the framing of the decision (as a potential gain or loss), students’ enjoyment of programming, perceived difficulty of programming, and their academic performance in the course. Our findings reveal intriguing insights into student risk-taking behavior. First, similar to software engineers in prior work, the framing of the decision significantly impacted the choices students made, with the loss framing leading to a higher likelihood for risky choices. Surprisingly, students displayed a greater inclination towards risk-taking compared to their professional counterparts in prior research. Furthermore, we observed that students’ prior academic performance in the course and their enjoyment of programming had a subtle influence on their risk-taking tendencies, with better-performing students and those who enjoyed programming being marginally more prone to taking risks. Notably, we did not find statistically significant correlations between perceived difficulty of programming and risk-taking behavior among students.
risk-taking behavior, student decision-making, framing effects, academic performance
$c$s: Social and professional topics Computing education
@@ 1. Introduction
Risk-taking is an inherent aspect of human decision-making, and part of daily life. From minor choices such as trying a new cuisine to more significant professional decisions such as how to best implement a new software feature, many factors contribute to how decisions are made. More complex decisions often involve carefully weighing up potential benefits against potential losses, taking into account an individual’s appetite for risk. Quite subtle factors can also play an important role, including how the choices that one is deciding between are presented. Indeed, the way that options are framed has been shown to significantly influence people’s choices, demonstrating that decision-making is as much about perception as it is about logic. In educational contexts, if the framing of decisions can impact the choices that students make, we believe this has important implications for educators.
While there exists a multitude of research on risk-taking in other fields (Yates, 1992), it has received little attention in computing education research. Our work is inspired by the very recent paper of Graf-Vlachy in the field of software engineering, published at ICSE in 2023  (Graf-Vlachy, 2023). Graf-Vlachy studied the risk-taking behavior of professional software engineers through a survey experiment with 124 participants. The primary finding was that the way decisions were presented (framed as gains or losses) significantly affected software engineers’ risk-taking behaviors. However, when adjusting for multiple testing, the study did not find a significant relationship between risk-taking and the Big Five personality traits. This prior work highlights that engineers are generally risk-averse, but framing can lead to substantial changes in their level of risk-taking. The results imply that both individual developers and project managers should be cognizant of framing effects when making decisions in software projects.
In our work, we partially replicate Graf-Vlachy’s study, but with students taking an introductory programming course. Firstly, we are interested in seeing if, similar to professionals, the framing of the risk-taking scenario correlates with students’ decisions. As the original study did not find significant correlations between personality traits and risk-taking, we analyze other attributes that might correlate with risk-taking behavior. More specifically, we examine the correlation between risk-taking and students’ performance in the course, their enjoyment of programming, and whether they find programming difficult. Lastly, we asked students to provide a rationale for their choice, and are interested in examining what rationales students give for their risk-taking choices. Our research questions for this work are as follows.
  • RQ1. Do framing and order of choices correlate with students’ risk-taking?
  • RQ2. What factors correlate with risk-taking?
  • RQ3. What rationales do students give for their risk-taking choices?
@@ 2.1. Exploring Framing Effects
This phenomenon – that a choice between a “sure thing” and a risky option of equal expected value is affected by option phrasing – has been widely replicated (Fagley and Miller, 1990; Kühberger, 1998; DeKay et al., 2022). A very recent study by Graf-Vlachy, published at the International Conference on Software Engineering (ICSE) in 2023, explored this in the domain of software engineering (Graf-Vlachy, 2023). Recognizing that risk-taking is inherent in daily software engineering decisions, Graf-Vlachy sought to provide empirical data on the subject, which is often overlooked in favor of risk management at the project or organizational level (Masso et al., 2022). A survey experiment was conducted with a sample of 124 software engineers. Participants were presented with a software engineering scenario and given two options: a risk-averse and a risk-seeking option. The framing manipulation involved presenting these options in two distinct ways: as “gains” and as “losses”.
The scenario asked participants to imagine they were working on a software project with a deadline, and that some requirements were wrongly implemented that will result in a missed deadline of 6 weeks. For participants in the first condition the options were framed as “gains”, i.e., there was a chance of recovering time. They were presented with the following two options: (A) If you reduce non-essential features, you will recover 2 weeks or (B) If you simplify the software architecture, there is a 1/3 chance that you will recover the full 6 weeks, and there is a 2/3 chance that the simplified architecture will lead to performance problems and you will not recover any time at all.
In the second condition, the options were described in terms of “losses”, i.e., the delay with which they would finish the project, and participants were asked to choose between: (A) If you reduce non-essential features, you will finish with a delay of 4 weeks or (B) If you simplify the software architecture, there is a 1/3 chance that you will finish the project with no delay at all, and there is a 2/3 chance that the simplified architecture will lead to performance problems and you will finish with a delay of 6 weeks.
These subtle differences in phrasing were used to study the framing effect on risk-taking decisions. No correlations were found between personality traits and risk-taking after multiple testing adjustments, suggesting that individual predispositions might not be as influential as situational framing. The key finding was the substantial effect of framing: software engineers exhibited higher risk-taking behaviors (i.e. favoring options involving an element of chance) when decisions were framed in terms of losses. The results imply that both individual developers and project managers should be cognizant of framing effects when making decisions in software projects.
@@ 2.2. Framing Effects in Education
To the best of our knowledge, there is no research on how framing effects might impact students in computing courses. We can see numerous scenarios where students might be faced with complex decisions as they are studying, and insights into the factors that might influence risk-taking behaviors for students can inform pedagogical strategies. In this work we seek to replicate the study described by Graf-Vlachy, exploring framing effects in the context of a scenario common to computing students – submitting assignments on time.
@@ 3. Methods
3.1. Course
Our data was collected from a large first-year programming course taught in the winter term of 2023 at a large public university in New Zealand. A total of 889 students were enrolled in the course that term, and all were invited to participate in our study for a very small amount of course credit (approximately 0.25% of the final grade).
The size of the course is relevant because a sufficiently large number of participants is needed to achieve the statistical power necessary for detecting an effect, if one truly exists, with minimal risk of encountering false positives or negatives. In the study we are replicating, Graf-Vlachy conduct a detailed power analysis to determine that 124 is the minimum sample size in order to detect a medium effect size with a desired power of 0.8 (Graf-Vlachy, 2023) and, in fact, their email-based recruitment resulted in exactly 124 participants. The much larger cohort in our study (853 students participated) allows us to account for additional variability in the data, and detect potentially smaller effect sizes.
@@ 3.1. Course
At the start of week 7, the mid-point of the course, students were beginning a new module and were asked to reflect on their prior experiences learning to program up to that point. Responses were collected online through the course learning management tool. The following prompt was shown: “Now that you have finished your first programming module, read the following statements and reflect on how much you agree with them – respond on a scale from Strongly disagree to Strongly agree”, followed by the two statements:
@@ 3.2. Experiment
A total of 853 students participated in the experiment. Students were randomly presented one of four scenario options. Two groups had the gain framing, and two had the loss framing for the scenario. Within each condition, the order of the safe and risky choice was randomly selected, leading to a total of four groups. For each condition, there was a “safe” option where the outcome was guaranteed, and a “risky” option where the outcome was based on chance, randomly leading to a more positive or negative outcome compared to the “safe” option. The expected value was exactly the same for all choices. The number of students in each experiment group is outlined in Table 1. The exact phrasing of the scenario is presented next.
@@ 3.4. Variables
Our study involves a total of six variables: (1) order of the questions, (2) framing of the scenario, (3) enjoyment of programming, (4) finding programming difficult, (5) performance in the course, and (6) choice of risk. The detailed description of the variables and their use is as follows.
  • (1) Order of the questions: The order in which the choices were presented to the students was random. This allows examining if the ordering of the choices affects which option is selected. Coded as 0 = risky first, 1 = safe first.
  • (2) Framing of the scenario: The scenario was framed as gains or losses, similar to the study by Graf-Vlachy. Coded as 0 = losses, 1 = gains.
  • (3) Enjoyment of programming: Halfway through the course, students were asked to express their level of agreement with the statement I find programming enjoyable. Likert-answer between 1 and 5 (1 = “Strongly disagree”, 2 = “Disagree”, 3 = “Neutral”, 4 = “Agree”, 5 = “Strongly agree”).
  • (4) Finding programming difficult: Halfway through the course, students were asked to express their level of agreement with the statement I find programming difficult. Likert-answer between 1 and 5 (1 = “Strongly disagree”, 2 = “Disagree”, 3 = “Neutral”, 4 = “Agree”, 5 = “Strongly agree”).
  • (5) Performance in the course: We use students’ performance in the mid-term examination. Score ranged from 0 to 48 (which was the maximum score available).
  • (6) Choice of risk: Whether students chose the risky (uncertain) outcome or the safe (guaranteed) outcome.
@@
Figure 1. Distributions of variables.
Figure 1 shows the distribution of student responses to the questions on enjoyment and difficulty of programming, and scores in the midterm examination.
@@ 3.5. Analysis
To answer RQ1, “Do framing and order of choices correlate with students’ risk-taking?”, we conduct a Barnard’s exact test [^1] Using the SciPy (Virtanen et al., 2020) implementation: https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.barnard_exact.html (Barnard, 1947) to analyze whether there are statistically significant differences between the gain and loss framing (i.e., whether the framing correlates with the choice), and separately if there are differences between the order of choices (i.e., whether the order in which the choices are presented correlates with the choice).
To answer RQ2, “What factors correlate with risk-taking?”, we follow the method of Graf-Vlachy’s original study (Graf-Vlachy, 2023), and run probit regression [^2] Using the statsmodels (Seabold and Perktold, 2010) Python package: https://www.statsmodels.org/dev/generated/statsmodels.discrete.discrete_model.Probit.html to analyze how the different variables (framing, order, enjoyment, difficulty, performance) correlate with the choice students make in the scenario.
@@
To answer RQ3, “What rationales do students give for their risk-taking choices?”, we randomly sampled 100 responses to the question “Please explain why you chose that option in a sentence or two”. Two researchers jointly coded the data for commonly occurring themes. One response could potentially include multiple themes.
@@ 4. Results
4.1. RQ1: Do Framing and Order of Choices Correlate With Students’ Risk-Taking?
@@
4.2. RQ2: What Factors Correlate With Risk-Taking?
@@
Table 4 shows the results of the probit regression. From the table, it can be seen that two factors do not statistically significantly correlate with the choice of risky or safe option: order (p-value = 0.656) and whether the student found programming difficult (p-value = 0.799). However, the other three factors show a statistically significant correlation with the choice: framing (p-value ¡ 0.001), enjoyment of programming (p-value = 0.017), and performance in the coursee (p-value = 0.009). The marginal effects of the variables on the dependent variable can be seen in the dy/dx column. For example, one point increase in performance (score in the midterm exam) will increase the likelihood of making the risky choice by 0.0055.
4.3. RQ3: What Rationales Do Students Give for Their Risk-Taking Choices?
@@ 5. Discussion
We were able to replicate earlier findings (Graf-Vlachy, 2023; Fagley and Miller, 1990; Kühberger, 1998; DeKay et al., 2022) that have suggested the framing of choices correlates with risk-taking behavior. Interestingly, compared to previous results derived from professional software engineers using a similar scenario (i.e. working on a software project), student participants in our study reported being more willing to take risks. Graf-Vlachy  (Graf-Vlachy, 2023) reported 7 out of 63 (11.1%) participants in the ‘gain’ condition and 19 out of 61 (31.1%) participants in the ‘loss’ condition chose the risky option. In contrast, in our case 190 out of 381 (49.9%) participants in the ‘gain’ condition and 290 out of 472 (61.4%) participants in the ‘loss’ condition chose the risky option. On the surface, this suggests that students might be more willing to take risks than professional software engineers, however this could also be explained by the fact that the stakes are lower in the context of course projects compared to professional contexts, where there may be impacts on employment or serious consequences for end users of production systems.
@@
Related to the factors that correlate with risk-taking behavior, we were surprised at the result that better performing students were more likely to take risks. Our initial hypothesis was that better performing students would be more similar to professional software engineers, i.e. less inclined to take risks. Conversely, poor performance may be explained by regular risk-taking involving studying and coursework. However, there are potential explanations for this. First, the study was conducted after the halfway point of the course, and we used the mid-term score as the measure of student performance. Students who did well in the midterm exam had therefore received positive feedback, and thus may have developed higher self-efficacy which may explain their confidence in taking risks to earn a higher mark. Better self-efficacy has been linked with better work-related performance (Stajkovic and Luthans, 1998), and is one of the strongest predictors of performance in introductory programming as well (Watson et al., 2014; Lishinski et al., 2016). We also found that students who enjoy programming more were more likely to take risks compared to those who do not. One possible explanation may be that such students are more likely to be excited by the prospect of modifying their current implementation (which was the “risky” option in our experiment).
@@
There has been some interest in computing education research on “nudging” strategies, where students are gently encouraged towards adopting better learning habits, often through various forms of messaging (Brown and Parnin, 2021; Zamprogno et al., 2020; Fouh et al., 2021; Edwards and Li, 2020). Given the quite large effect we have observed in our work, which simply involved framing options differently, future work should explore how better framing of ‘nudges’ could lead to more impactful results.

### v3

@@ Abstract.
Understanding how individuals make decisions involving risk is a fundamental aspect of behavioral research. Despite the ubiquity of risk in various aspects of life, limited empirical work has explored student risk-taking behavior in computing education. This study aims to partially replicate prior research on risk-taking behavior in software engineers while focusing on students, shedding light on the factors that affect their risk-taking choices. In our work, students were presented with a hypothetical scenario related to meeting a course project deadline, where they had to choose between a risky option and a safer alternative. We examined several factors that might influence these choices, including the framing of the decision (as a potential gain or loss), students’ enjoyment of programming, perceived difficulty of programming, and their performance in the course. Our findings reveal intriguing insights into student risk-taking behavior. Similar to software engineers in prior work, the framing of the decision significantly impacted the choices students made, with the loss framing leading to a higher likelihood for risky choices. Surprisingly, students displayed a greater inclination towards risk-taking compared to their professional counterparts in prior research. Furthermore, we observed that students’ performance in the course and their enjoyment of programming had a subtle correlation with their risk-taking tendencies, with better-performing students and those who enjoyed programming being marginally more prone to taking risks. Notably, we did not find statistically significant correlations between perceived difficulty of programming and risk-taking behavior among students.
risk-taking behavior, student decision-making, framing effects, academic performance, replication, nudging
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 2024 ACM Virtual Global Computing Education Conference V. 1; December 5–8, 2024; Virtual Event, NC, USAbooktitle: Proceedings of the 2024 ACM Virtual Global Computing Education Conference V. 1 (SIGCSE Virtual 2024), December 5–8, 2024, Virtual Event, NC, USAdoi: 10.1145/3649165.3690097isbn: 979-8-4007-0598-4/24/12ccs: Social and professional topics Computing education
@@ 1. Introduction
Risk-taking is an inherent aspect of human decision-making, and part of daily life. From minor choices such as trying a new cuisine to more significant professional decisions such as how to best implement a new software feature, many factors contribute to how decisions are made. More complex decisions often involve carefully weighing up potential benefits against potential losses, taking into account an individual’s appetite for risk. Quite subtle factors can also play an important role, including how the choices that one is deciding between are presented. Indeed, the way that options are framed has been shown to significantly influence people’s choices (Tversky and Kahneman, 1981), demonstrating that decision-making is as much about perception as it is about logic. In educational contexts, if the framing of decisions can impact the choices that students make, we believe this has important implications for educators.
While there exists a multitude of research on risk-taking in other fields (Yates, 1992), it has received little attention in computing education research. Our work is inspired by the work of Graf-Vlachy in the field of software engineering (Graf-Vlachy, 2023) where the risk-taking behavior of professional software engineers was studied through a survey experiment with 124 participants. The primary finding was that the way decisions were presented (framed as gains or losses) significantly correlated with software engineers’ risk-taking behaviors. However, when adjusting for multiple testing, the study did not find a significant relationship between risk-taking and the Big Five personality traits (see (Roccas et al., 2002)). This prior work highlights that engineers are generally risk-averse, but framing can lead to substantial changes in their level of risk-taking. The results imply that both individual developers and project managers should be cognizant of framing effects when making decisions in software projects.
In our work, we partially replicate Graf-Vlachy’s study with students taking an introductory programming course. We examine if, similar to professionals, the framing of the risk-taking scenario correlates with students’ decisions. As the original study did not find significant correlations between personality traits and risk-taking, we analyze other attributes. More specifically, we study the correlation between risk-taking and students’ performance in the course, their enjoyment of programming, and whether they find programming difficult. We also explore students’ rationales for their risk-taking choices. Our research questions for this work are:
  • RQ1. How does the framing of a risk-taking scenario, the order of choices, the perceived difficulty of programming and programming enjoyment, and course performance correlate with risk-taking?
  • RQ2. What rationales do students give for their risk-taking choices?
@@ 2.1. Exploring Framing Effects
This phenomenon – that a choice between a “sure thing” and a risky option of equal expected value is affected by option phrasing – has been widely replicated (Fagley and Miller, 1990; Kühberger, 1998; DeKay et al., 2022). Graf-Vlachy (Graf-Vlachy, 2023) explored this in the domain of software engineering, recognizing that risk-taking is inherent in daily software engineering decisions. His study sought to provide empirical data on the subject, which is often overlooked in favor of risk management at the project or organizational level (Masso et al., 2022). Professional software developers were given a survey with a software engineering scenario and given two options: a risk-averse and a risk-seeking option. The framing manipulation involved presenting these options in two distinct ways: as ‘gains’ and as ‘losses’.
The scenario asked participants to imagine they were working on a software project with a deadline, and that some requirements were wrongly implemented which will result in a missed deadline of 6 weeks. For participants in the first condition the options were framed as ‘gains’, i.e., there was a chance of recovering time. They were presented with the following two options: (A) If you reduce non-essential features, you will recover 2 weeks or (B) If you simplify the software architecture, there is a 1/3 chance that you will recover the full 6 weeks, and there is a 2/3 chance that the simplified architecture will lead to performance problems and you will not recover any time at all.
In the second condition, the options were described in terms of ‘losses’, i.e., the delay with which they would finish the project, and participants were asked to choose between: (A) If you reduce non-essential features, you will finish with a delay of 4 weeks or (B) If you simplify the software architecture, there is a 1/3 chance that you will finish the project with no delay at all, and there is a 2/3 chance that the simplified architecture will lead to performance problems and you will finish with a delay of 6 weeks.
These subtle differences in phrasing were used to study the framing effect on risk-taking decisions. No correlations were found between personality traits and risk-taking after multiple testing adjustments, suggesting that individual predispositions might not be as influential as situational framing. The key finding was the substantial effect of framing: software engineers exhibited higher risk-taking behaviors (i.e., favoring options involving an element of chance) when decisions were framed in terms of losses. The results imply that both individual developers and project managers should be cognizant of framing effects when making decisions in software projects.
@@ 2.2. Framing Effects in Education
To the best of our knowledge, there is no research on how framing effects might impact students in computing courses. However, students are faced with complex decisions as they are studying, and insights into the factors that might influence risk-taking behaviors for students can inform pedagogical strategies. We seek to replicate the study of Graf-Vlachy, exploring framing effects in the context of a scenario common to computing students – submitting assignments on time (Denny et al., 2021; Castro et al., 2022; Leinonen et al., 2021, 2022).
@@ 3. Methods
3.1. Context
Our data was collected from a large first-year programming course taught in 2023 at a large public university in New Zealand. A total of 889 students were enrolled in the course that term, and all were invited to participate in our study for a very small amount of course credit (approximately 0.25% of the final grade).
The size of the course is relevant because a sufficiently large number of participants is needed to achieve the statistical power necessary for detecting an effect, if one truly exists, with minimal risk of encountering false positives or negatives. In the study we are replicating, Graf-Vlachy conduct a detailed power analysis to determine that 124 is the minimum sample size in order to detect a medium effect size with a desired power of 0.8 (Graf-Vlachy, 2023). The much larger cohort in our study (853 students participated) allows us to account for additional variability in the data, and detect potentially smaller effect sizes.
@@ 3.1. Context
At the start of week 7, the mid-point of the course, students were beginning a new module and were asked to reflect on their prior experiences learning to program up to that point. Responses were collected online through the course learning management system. The following prompt was shown: “Now that you have finished your first programming module, read the following statements and reflect on how much you agree with them – respond on a scale from Strongly disagree to Strongly agree”, followed by the two statements:
@@ 3.2. Experiment
The study had ethics approval from the university’s human participants ethics committee. Students were randomly presented one of four scenario options. Two groups had the gain framing, and two had the loss framing for the scenario. Within each condition, the order of the safe and risky choice was randomly selected, leading to a total of four groups. For each condition, there was a “safe” option where the outcome was guaranteed, and a “risky” option where the outcome was based on chance, randomly leading to a more positive or negative outcome compared to the “safe” option. The expected value was exactly the same for all choices. The number of students in each experiment group is outlined in Table 1.
@@ 3.4. Variables
Our study involves a total of six variables. The detailed description of the variables and their use is as follows.
  • Order of the questions: The order in which the choices were presented to the students was random. This allows examining if the ordering of the choices affects which option is selected. Coded as 0 = risky first, 1 = safe first.
  • Framing of the scenario: The scenario was framed as gains or losses, similar to the study by Graf-Vlachy. Coded as 0 = losses, 1 = gains.
  • Enjoyment of programming: Halfway through the course, students were asked to express their level of agreement with the statement I find programming enjoyable. Likert-answer between 1 and 5 (1 = “Strongly disagree”, 2 = “Disagree”, 3 = “Neutral”, 4 = “Agree”, 5 = “Strongly agree”).
  • Finding programming difficult: Halfway through the course, students were asked to express their level of agreement with the statement I find programming difficult. Likert-answer between 1 and 5 (1 = “Strongly disagree”, 2 = “Disagree”, 3 = “Neutral”, 4 = “Agree”, 5 = “Strongly agree”).
  • Performance in the course: We use students’ performance in the mid-term examination. Score ranged from 0 to 48 (which was the maximum score available).
  • Choice of risk: Whether students chose the risky (uncertain) outcome or the safe (guaranteed) outcome. Coded as 0 = safe, 1 = risky.
@@
Figure 1. Distributions of variables. For (a) and (b), 1 = “Strongly disagree”, 2 = “Disagree”, 3 = “Neutral”, 4 = “Agree”, and 5 = “Strongly agree”.
@@ 3.5. Analysis
To answer RQ1, “How does the framing of a risk-taking scenario, the order of choices, the perceived difficulty of programming and programming enjoyment, and course performance correlate with risk-taking?”, we first conduct a Barnard’s exact test [^1] Using the SciPy (Virtanen et al., 2020) implementation: https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.barnard_exact.html (Barnard, 1947) to analyze whether there are statistically significant differences between the gain and loss framing (i.e., whether the framing correlates with the choice), and separately if there are differences between the order of choices (i.e., whether the order in which the choices are presented correlates with the choice).
Then, we follow the method of Graf-Vlachy’s original study (Graf-Vlachy, 2023), and run probit regression [^2] Using the statsmodels (Seabold and Perktold, 2010) Python package: https://www.statsmodels.org/dev/generated/statsmodels.discrete.discrete_model.Probit.html to analyze how the different variables (framing, order, enjoyment, difficulty, performance) correlate with the choice students make in the scenario.
@@
To answer RQ2, “What rationales do students give for their risk-taking choices?”, we randomly sampled 100 responses to the question “Please explain why you chose that option in a sentence or two”. Two researchers jointly coded the data for commonly occurring themes, discussing and resolving any disagreements. One response could include multiple themes.
@@ 4. Results
4.1. RQ1: How Do Different Factors Correlate With Risk-Taking?
@@
Table 4 shows the results of the probit regression. Two factors do not statistically significantly correlate with the choice of risky or safe option: order (p-value = 0.656) and whether the student found programming difficult (p-value = 0.799). However, the other three factors show a statistically significant correlation with the choice: framing (p-value ¡ 0.001), enjoyment of programming (p-value = 0.017), and performance in the course (p-value = 0.009). The marginal effects of the variables on the dependent variable can be seen in the dy/dx column. For example, one point increase in performance (score in the midterm exam) will increase the likelihood of making the risky choice by 0.0055.
Figure 1 shows the distribution of student responses to the questions on enjoyment and difficulty of programming, and scores in the midterm examination. The distribution of midterm scores is similar to previous iterations of the course.
4.2. RQ2: What Rationales Do Students Give for Their Risk-Taking Choices?
@@ 5. Discussion
We were able to replicate earlier findings (Graf-Vlachy, 2023; Fagley and Miller, 1990; Kühberger, 1998; DeKay et al., 2022) that have suggested the framing of choices correlates with risk-taking behavior. Interestingly, compared to previous results derived from professional software engineers using a similar scenario (i.e., working on a software project), student participants in our study reported being more willing to take risks. Graf-Vlachy  (Graf-Vlachy, 2023) reported 7 out of 63 (11.1%) participants in the ‘gain’ condition and 19 out of 61 (31.1%) participants in the ‘loss’ condition chose the risky option. In contrast, in our case 190 out of 381 (49.9%) participants in the ‘gain’ condition and 290 out of 472 (61.4%) participants in the ‘loss’ condition chose the risky option. On the surface, this suggests that students might be more willing to take risks than professional software engineers, however this could also be explained by the fact that the stakes are lower in the context of course projects compared to professional contexts, where there may be impacts on employment or serious consequences for end users of production systems.
@@
Related to the factors that correlate with risk-taking behavior, we were surprised at the result that better performing students were more likely to take risks. Our initial hypothesis was that better performing students would be more similar to professional software engineers, i.e., less inclined to take risks. Conversely, poor performance may be explained by regular risk-taking involving studying and coursework. However, there are potential explanations for this. First, the study was conducted after the halfway point of the course, and we used the mid-term score as the measure of student performance. Students who did well in the midterm exam had therefore received positive feedback, and thus may have developed higher self-efficacy which may explain their confidence in taking risks to earn a higher mark. Better self-efficacy has been linked with better work-related performance (Stajkovic and Luthans, 1998), and is one of the strongest predictors of performance in introductory programming as well (Watson et al., 2014; Lishinski et al., 2016). We also found that students who enjoy programming more were more likely to take risks compared to those who do not. One possible explanation may be that such students are more likely to be excited by the prospect of modifying their current implementation (which was the “risky” option in our experiment).
@@
The results of our study have implications for computing instructors. There has been some interest in computing education on “nudging” strategies, where students are gently encouraged towards adopting better learning habits, often through various forms of messaging (Brown and Parnin, 2021; Zamprogno et al., 2020; Fouh et al., 2021; Edwards and Li, 2020; Ilves et al., 2018; Indriasari et al., 2023; Rogers et al., 2021). Given the quite large effect we have observed in our work, which simply involved framing options differently, future work should explore how better framing of ‘nudges’ could lead to more impactful results. Another area that could utilize our results is assessment design. Instructors could strategically leverage framing in how assessments are presented. If the goal is to encourage experimentation and creative problem-solving, a loss-framed approach might be effective, as it promotes risk-taking. Conversely, if the focus is on reinforcing foundational skills or ensuring consistent performance, positive framing could help guide students toward safer, more measured choices.